# SPIN-Bench: How Well Do LLMs Plan Strategically and Reason Socially?

### v3

@@ 5 GPT-4o Austria, England
    Support-Self Orders: Support-Self orders are the orders the agent issue to make its unit to support its own unit’s attacking other power’s unit. We have three metrics for the support-self orders issued by the agent: Successful support-self count, Total support-self count, and Successful support-self rate. This metric evaluates model’s multi-step reasoning ability and the planning-action consistency. It is the key to a successful attack order.
@@
    Support-Others Orders: Support-Other orders are the orders the agent issue to make its unit to support other’s unit’s attacking another power’s unit. Supporting others means helping others, which is a highly strategic action in a such competitive game setting. We have three metrics for the support-others orders issued by the agent: Successful support-others count, Total support-others count, and Successful support-others rate. This metric evaluates model’s social intelligence, theory of mind, and its strategic reasoning on building breakable but profitable relationships with its opponents.
@@
Acceptance Rate=Number of Accepted ProposalsTotal $\text{Acceptance Rate}\;=\;\frac{\text{Number of Accepted Proposals}}{\text{Total Proposals}}$
@@
(5) Perspective Taking or Empathy: Here we label whether the message demonstrates an explicit acknowledgment of the other party’s viewpoint or needs. We mark 1 if the agent references or empathizes with the recipient’s perspective (e.g., “I understand you’re concerned about Italy in the Mediterranean”), or 0 if it does not. High values of perspective taking can show advanced “theory of mind” capabilities. In the metric, we compute the ratio of the perspective taking messages to the total number of messages.
@@
For each negotiation phase in a complete game trajectory (20 years), we provide the annotation prompt to o1 model, gathering all the results from its analysis, and compute the statistics. An example of one negotiation phase and o1 model’s annotation is shown in Appendix D.3.
@@ Appendix C Supplementary Experiments
C.1 Two-player Competitive Games
@@
LLM Tic Tac Toe (%$\uparrow$ Connect Four (%$\uparrow$ Chess (%$\uparrow$
@@
C.2 Diplomacy
C.2.1 Basic Skill Evaluation
@@
C.2.2 Multi-agent Experiment
@@
o1 87/48 116/57 0.75/0.84 21/9 34/12 0.62/0.75 27/22 62/47 0.44/0.47 0/1 0/1 0.00/1.00 15/4 36/8
GPT-4o 17/6 47/10 0.36/0.60 5/1 27/4 0.19/0.25 4/1 14/7 0.29/0.14 0/0 0/1 0.00/0.00 1/6 3/8
GPT-4-turbo 26/18 61/28 0.43/0.64 9/5 35/12 0.26/0.42 15/8 77/31 0.19/0.26 2/0 5/3 0.40/0.00 0/1 3/3
Table 10: Performance indicators for three models (o1, GPT-4o, GPT-4-turbo) across various agent settings in Diplomacy (2/3/4/5 agents), presented as without negotiation / with negotiation (x/y). Each row shows: (Move Order) success number / total number / success rate, (Attack Order) success number / total number / success rate, (Support-Self Order) success number / total number / success rate, (Support-Others Order) success number / total number / success rate, then the number of supply centers(SC), and controlled regions(CR) at the end of game.
C.2.3 Human Validation of LLM Annotations
To validate the reliability of our LLM-assisted metrics, we conducted a human evaluation experiment to assess the accuracy of the annotations generated by Claude 3.7 Sonnet. Our research team manually annotated 5% of all negotiation messages across all six metrics. We then calculated the agreement rate between human annotations and LLM annotations to measure the reliability of our automated evaluation approach.
Table 11 presents the agreement rates for each negotiation metric. The results demonstrate high agreement between human and LLM annotations across all metrics, with accuracy rates ranging from 91.67% to 100%. This strong agreement suggests that the LLM-based annotation approach provides reliable evaluations for analyzing negotiation behaviors in Diplomacy.
@@ C.2.3 Human Validation of LLM Annotations
Appendix D Prompt Design and Interaction Flow
D.1 PDDL
@@ 5. Solution Output:
D.2 Cooperative Game: Hanabi
@@
- Hidden info: XX’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: XX’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: XX’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: XX’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: XX’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: XX’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: XX’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: XX’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: XX’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: XX’. No hints about this cards color or rank have been given yet.
@@
D.3 Strategic Game: Diplomacy
@@
1.1. Recap each message from the last round, identifying who said what.
@@
2.2. Summarize your opponents positions (who seems strong, who seems weak, who might be desperate).
@@ 3.2. Whether you should move into other powers unit to attack?
3.3. If you dont have enough unit to attack some region, do you want to move your other unit to the adjacent region to prepare for the attack in the next round?
@@ 5.2. You should take control more and more supply centers
5.5. IIterate over your planned attack targets, and compute how many other powers units are in that target location. Next, determine how many of your own units and how much support you need to succeed in the attack. You cannot attack with fewer supports than the defending unit. Verify that the total of your supports plus one is greater than the defending units strength. If this condition is not met, return to the beginning and think again.

### v4

@@ 5 GPT-4o Austria, England
Table 6: Agent Assignments for Multi-Agent Experiment Settings
@@
    Support-Self Orders(SS): Support-Self orders are the orders the agent issue to make its unit to support its own unit’s attacking other power’s unit. We have three metrics for the support-self orders issued by the agent: Successful support-self count, Total support-self count, and Successful support-self rate. This metric evaluates model’s multi-step reasoning ability and the planning-action consistency. It is the key to a successful attack order.
@@
    Support-Others Orders(SO): Support-Other orders are the orders the agent issue to make its unit to support other’s unit’s attacking another power’s unit. Supporting others means helping others, which is a highly strategic action in a such competitive game setting. We have three metrics for the support-others orders issued by the agent: Successful support-others count, Total support-others count, and Successful support-others rate. This metric evaluates model’s social intelligence, theory of mind, and its strategic reasoning on building breakable but profitable relationships with its opponents.
@@
$\text{Acceptance Rate}\;=\;\frac{\text{Number of Accepted Proposals}}{\text{Total Proposals}}$
@@
(5) Perspective Taking or Empathy: Here we label whether the message demonstrates an explicit acknowledgment of the other party’s viewpoint or needs. We mark 1 if the agent references or empathizes with the recipient’s perspective (e.g., “I understand you’re concerned about Italy in the Mediterranean”), or 0 if it does not. High values of perspective taking can show advanced “theory of mind” capabilities. It is an unambiguous, observable signal that the agent is actively modeling the other party’s mental state. Explicit perspective-taking often leads to more persuasive proposals in negotiation, which reflects a deeper level of social intelligence. In the metric, we compute the ratio of the perspective taking messages to the total number of messages.
@@
For each negotiation phase in a complete game trajectory (20 years), we provide the annotation prompt to Claude 3.7 Sonnet model, gathering all the results from its analysis, and compute the statistics. An example of how to annotate one message and LLM’s annotation is shown in Appendix E.3.
@@ Appendix C Supplementary experiments
C.1 PDDL
Figure 6: Evaluation of LLM performance in retrieving specific states from full-information trajectories. Each dot indicates the average accuracy for an individual task setting, computed over trajectories with lengths ranging from 1 to 50.
Factual Retrieval in Classical Planning.
Here, we investigate whether LLMs can reliably retrieve key facts from a planning trajectory. For each task, we provide the model with the initial state ($s_{\text{ini}}$), the goal state ($S_{G}$), and trajectories that explicitly detail the actions along with their corresponding state transitions. We control the number of mutable properties; for example, in a task involving moving balls, the ball location is considered mutable, whereas ball color might also be a property—but if no action can alter this property, we do not count it as mutable. We then prompt the model to report the state at a specified step. Figure 6 illustrates how retrieval accuracy varies with trajectory length. Notably, o1 performs most consistently, confirming that it ”reads” multi-step expansions more accurately than either GPT-4o or o1-mini.
C.2 Two-player competitive games
@@
[…]
@@
C.3 Diplomacy
C.3.1 Basic skill evaluation
@@
C.3.2 Multi-agent experiment
@@
o1 87/48 116/57 21/9 34/12 27/22 62/47 0/1 0/1 15/4 36/8
GPT-4o 17/6 47/10 5/1 27/4 4/1 14/7 0/0 0/1 1/6 3/8
GPT-4-turbo 26/18 61/28 9/5 35/12 15/8 77/31 2/0 5/3 0/1 3/3
Table 10: Performance indicators for three models (o1, GPT-4o, GPT-4-turbo) across various agent settings in Diplomacy (2/3/4/5 agents), presented as without negotiation / with negotiation (x/y). Each row shows: (Move Order) success number / total number, (Attack Order) success number / total number, (Support-Self Order) success number / total number, (Support-Others Order) success number / total number, then the number of supply centers(SC), and controlled regions(CR) at the end of game. Success rates are omitted to save space.
C.3.3 Human Validation of LLM Annotations
To validate the reliability of our LLM-assisted metrics, we conducted a rigorous human evaluation experiment. We randomly sampled 5% of all negotiation messages from our Diplomacy dataset and employed a message-level annotation approach designed to simplify the annotation process. Three trained human annotators, all members of our research team, independently labeled each sampled message across six binary classification metrics (defined in Appendix B.5.2). To ensure consistency between human and LLM annotations, annotators were provided with the same contextual information utilized by the LLM, including negotiation message content, identities of sender and recipient, current game state (map locations of supply centers, units, and controlled regions), all powers’ issued orders from the preceding negotiation phase, and the message sender’s private negotiation strategy.
Appendix D Discussion/limitation
D.1 Limited complixty on ”social Intelligence”
Social intelligence encompasses a broad set of abilities, spanning both social awareness social facility. Within this framework, SPIN-Bench is intentionally scoped to address one key aspect from each of Goleman’s broad categories: Theory of Mind, which aligns with ’social awareness’, and negotiation, which falls under ’social facility’.
D.2 Theory of Mind: Explicit vs. Implicit Perspective-Taking
Theory of Mind capabilities represent a critical aspect of social intelligence. In our study, we mainly examined explicity ToM. For each metric, we used a simple majority vote among annotators to derive the ground-truth human labels. Inter-annotator agreement, measured as the percentage of cases where all three annotators assigned identical binary labels, was 87%, indicating robust consistency for this classification task. Subsequently, we compared these majority-vote human labels to annotations generated by Claude 3.7 Sonnet. The agreement between Claude’s annotations and human annotators ranged from 91.67% to 100% across different metrics, demonstrating strong concordance. Table 11 provides a detailed breakdown of these agreement rates. This high level of alignment between human and LLM annotations affirms the reliability and effectiveness of our LLM-based annotation methodology for evaluating negotiation behaviors in Diplomacy.
@@ D.1 Limited complixty on ”social Intelligence”
Appendix E Prompt design and interaction flow
E.1 PDDL
@@ 5. Solution Output:
E.2 Cooperative game: Hanabi
@@
- Hidden info: $X$’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: $X$’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: $X$’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: $X$’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: $X$’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: $X$’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: $X$’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: $X$’. No hints about this cards color or rank have been given yet.
@@
- Hidden info: $X$’. This represents what you cannot see about this card. It means you have no direct knowledge about the cards identity from your perspective.
- Known info: $X$’. No hints about this cards color or rank have been given yet.
@@
E.3 Strategic game: Diplomacy
@@
$1.$ Recap each message from the last round, identifying who said what.
@@
$2.$ Summarize your opponents positions (who seems strong, who seems weak, who might be desperate).
@@ 3.2. Whether you should move into other powers unit to attack?
$3.$ If you dont have enough unit to attack some region, do you want to move your other unit to the adjacent region to prepare for the attack in the next round?
@@ 5.2. You should take control more and more supply centers
$5.$ IIterate over your planned attack targets, and compute how many other powers units are in that target location. Next, determine how many of your own units and how much support you need to succeed in the attack. You cannot attack with fewer supports than the defending unit. Verify that the total of your supports plus one is greater than the defending units strength. If this condition is not met, return to the beginning and think again.